OpenStudy (anonymous):

is my answer: 2x^3 - x^ -13x -15 to: (x)= 2x^3-5x; g(x)=x^2+8x+15

OpenStudy (isaiah.feynman):

What?

OpenStudy (anonymous):

for Find f − g and its domain.

OpenStudy (jdoe0001):

\(\bf f(x)= 2x^3-5\qquad \qquad g(x)=x^2+8x+15\\ \quad \\ (f-g)(x) =[2x^3-5]\quad -\quad [x^2+8x+15] \)

OpenStudy (anonymous):

i don't understand. my answer for:Find f + g and its domain. would be 2x^3 + x^2 + 3x + 15 so would that mean if I am looking for f-g it would be: -2x^3 - x^2 -13x -15

OpenStudy (jdoe0001):

\(\bf f(x)= 2x^3-5\qquad \qquad g(x)=x^2+8x+15\\ \quad \\ (f-g)(x) =[2x^3-5] - [x^2+8x+15]\implies 2x^3-5-x^2-8x-15\\ \quad \\ 2x^3-x^2-8x-20\)

OpenStudy (anonymous):

what do you mean by qquad?